# On the Adversarial Robustness of Learning-based Conformal Novelty Detection

### v3

@@ III-C Surrogate Decision-based Attack Scheme
Motivated by our oracle setting, we consider a practical scenario, where the attacker does not have the training samples $\{Z_{j}\}_{j=1}^{n}$ and does not know the true labels of the test samples, but is allowed to query the label from the user who applies AdaDetect, with underlying machine learning algorithms unknown to the attacker. Such query access is a standard assumption in adversarial settings as mentioned in Section I-B, reflecting a realistic constraint on the attacker’s capability. Specifically, the attack has
@@
Query: The attacker can query the user (who owns all the training and test data and the AdaDetect algorithm) to obtain the labels for all the testing data $\{Z_{j}\}_{j=n+1}^{m+n}$.
We propose a surrogate score function $g(z)$, trained on the pseudo-labeled dataset $\mathcal{D}=\{(Z_{n+i},y_{i})\}_{i=1}^{m}$, where we refer to $\{y_{i}\}_{i=1}^{m}$ as pseudo-labels since , unlike the true labels available in our oracle setting, these are the labels assigned by AdaDetect on the entire test set. This surrogate score function approximates AdaDetect’s binary decision: $\mathbf{1}\{g(Z_{n+i})\geq 0.5\}\approx y_{i}$, enabling decision-based adversarial attacks on a black-box detector.
@@
Step 2: Surrogate score function training. Assuming AdaDetect has reasonable detection power, we form the pseudo-labeled dataset $\newline \mathcal{D}=\{(Z_{n+i},Y_{i})\}_{i=1}^{m}$ and train a surrogate score function
$g(z)\leftarrow\mathrm{TrainScoreFunction}(\mathcal{D}).$
@@
$\widetilde{Z}_{n+i}=f_{\text{attack}}(Z_{n+i},g(z)).$
It is important to note that our surrogate decision-based attack does not require information about the algorithm, making it a practical attack scheme. This point is also highlighted in our experiment section through mismatched setups where the user and attacker adopt two different algorithms to learn the score function (see Experiment 3 and Experiment A.1 for details).
@@ IV Experiments
As explained in Section III-A, we consider both the Adadetect scheme and the one by Bates el al. [4], using two adversarial machine learning attacks: HSJA and Boundary attack.
@@ IV-A Applying Surrogate Scheme on Adadetect
We evaluated the attack performance on three types of synthetic data distributions: independent Gaussian, non-Gaussian, and exchangeable Gaussian data, and four real-world datasets (Shuttle, Credit Card, KDDCup99, and Mammography) with diverse characteristics and application domains (see data descriptions in Appendix F). All experiments use the following base parameters: training sample size $n=5000$, testing sample size $m=1000$, $k=4000$, true null data $m_{0}=900$, and significance level $\alpha=0.1$. Each experiment is repeated $20$ times to calculate the mean and variance of the FDR and power.
Due to space limitations, we defer the synthetic data generation and experiment results to Appendix F, along with one more real-world data experiment.
Experiment 1: Real-world data with RF models. We evaluated both oracle and surrogate attack performance in the four real-world datasets using identical RF architectures for both score functions, with attack size $m_{a}=200$. This allows us to compare attack effectiveness across different real-world data characteristics. We compare the boundary attack with our default HSJA under both oracle and surrogate attack schemes, using identical RF architectures for both score functions with $m_{a}=200$. The result shows that our oracle and surrogate attack schemes and decision-based algorithms (HopSkipJump and Boundary) can significantly increase the FDR in comparison to the original FDR.
@@
TABLE II: Experiment 1: Power + RF
@@
We make two important observations. Firstly, compared to the other three datasets, the original power on Mammography is relatively low ($\sim 0.48$). As a consequence, the surrogate method learns the score function from less accurate labels, leading to a larger gap in FDR between the oracle vs. the surrogate. Similar phenomena appear in the next real-world experiments, as well as the two in the Appendix F. Secondly, the power after attack often increases, since the attack targets data points near the decision boundary in practice, some of which belong to the alternative hypothesis and are thus more likely to be correctly rejected.
Experiment 2: Real-world data with mismatched configurations. We apply mismatched score function configurations (RF–NN) to the four real-world datasets with a fixed attack size of $m_{a}=200$, evaluating both the oracle and surrogate attack performance. This enables us to assess how different model combinations affect each attack type’s performance across various real-world scenarios. The results indicate that an attacker can employ a model different from the user’s and still inflate the FDR beyond the target level. Our experiments show that using the neural network configuration as the attacker’s model can substantially speed up the attack process. The results for RF-RF configuration is covered in Experiment $2$.
TABLE III: Experiment 2: FDR + RF-NN
@@
TABLE IV: Experiment 3 with Monte Carlo adjustment
@@
TABLE V: Experiment 3 without Monte Carlo adjustment
@@
To demonstrate the generality of our attack framework, we evaluate the adversarial robustness of the one-class classifier-based method proposed by Bates et al. in [4]. This approach serves as a fundamental baseline. We utilize two real-world datasets and a synthetic Gaussian data distribution. The null (e.g., non-fraud) data is partitioned into a training set of size $n_{\text{train}}=4000$ and a calibration set of size $n_{\text{cal}}=1000$. The test set $\mathcal{D}_{\text{test}}$ consists of $m=2000$ samples, comprising $m_{0}=1800$ inliers (normal transactions) and $m_{1}=200$ outliers (e.g., fraud). The p-values are computed using one-class classifier-based scheme, and the BH procedure is applied for FDR control.
@@
(2) Surrogate Training with Label Flipping: The attacker queries the one-class classifier-based scheme to obtain initial binary rejections. To train the surrogate model $g(z)$ (a Multi-Layer Perceptron with one hidden layer of 100 units), we employ a heuristic label-flipping strategy: $75\%$ of the rejected samples with the highest p-values are relabeled as inliers in the training set. This encourages the surrogate to learn a more restrictive decision boundary, as only the most extreme outliers retain their original labels. Since HSJA and Boundary attack are decision-based attack, this surrogate model with tighter decision boundary for outliers will encourage those two decision-based attack to be more aggressive. We remark that a more conservative boundary (e.g., using a $50\%$ threshold or no flip at all) often fails here because it results in a loose rejection region, making it difficult for the attack to successfully push a sample labeled inlier into the real outlier space. The sample may successfully move into this loose boundary for outliers, but it’s still not extreme enough for one-class classifier-based scheme to notice. While the optimal threshold may vary across different datasets, this parameter serves as a tuning knob: increasing the flipping percentage makes the HSJA and Boundary attack more aggressive, facilitating more effective adversarial generation.
(3) Adversarial Generation: We utilize HSJA and Boundary attack on the surrogate model $g(z)$ to generate adversarial perturbations for the target set $\mathcal{A}$. The attack is aiming to cross the decision boundary.
Experiment 3: Both real-world and synthetic data on Bates’s method. We evaluated the attack performance at significance levels $\alpha=0.1$ both with and without Monte Carlo adjustment (recall Remark 1). The results quantify whether our surrogate approach succeed at attacking the standard one-class classifier-based scheme.
@@ V Discussion
Appendix A Proof of Lemma 1
Proof.
and we write this in a compact form as $\widetilde{U}_{\mathcal{A}}=f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)$. We want to show that
According to Proposition 1, $f_{\text{attack}}$ does not depend on order of the elements from $U_{k+1}$ to $U_{n+\tilde{m}_{0}}$ in $U_{\setminus\mathcal{A}}$, so we have $\widetilde{U}_{\mathcal{A}}=f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}}^{\pi},U_{\mathcal{A}}\cup V)$. For any measurable set $\mathcal{U}_{\setminus\mathcal{A}}$ in the support of $U_{\setminus\mathcal{A}}$, we have
$\mathrm{P}(U_{\setminus\mathcal{A}}\in\mathcal{U}_{\setminus\mathcal{A}},f_{\text{attack}}(U_{\mathcal{A}};U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}}|V=v)=\mathrm{P}(U_{\setminus\mathcal{A}}^{\pi}\in\mathcal{U}_{\setminus\mathcal{A}},f_{\text{attack}}(U_{\mathcal{A}};U_{\setminus\mathcal{A}}^{\pi},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}}|V=v).$
Therefore,
Appendix B Proof of Lemma 2
Proof.
Let
$Q=h(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V)=((Z_{1},\dots,Z_{k}),\{Z_{k+1},...,Z_{n},Z_{n+i}:i\in\mathcal{H}_{1}\cup\mathcal{H}_{0}\setminus A\}.$ (14)
By Lemma 1, we have
This implies
Since the permutation keeps the first $k$ indices fixed, by the definition of $Q$, we have $h(U_{\setminus\mathcal{A}}^{\pi},\widetilde{U}_{\mathcal{A}},V)=h(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V)=Q$. Thus
Observe that the score terms $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$ are deterministic functions of the conditional variables $\widetilde{U}_{\mathcal{A}},V,Q$. Since $\pi(i)=i$ for all $i\leq k$, the exchangeability is preserved as follows,
$(S_{k+1},\dots,S_{n},S_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})\overset{d}{=}$
$(S_{\pi(k+1)},\dots,S_{\pi(n)},S_{\pi(n+i)}:i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A}).$
Appendix C Proof of Theorem 1
@@
We can decompose $\text{FDR}_{\text{attack}}$ as follows,
$\hskip-1.00006pt\text{FDR}_{\text{attack}}\hskip-1.99997pt=\hskip-1.99997pt\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}\right]+\mathbb{E}\left[\sum_{i\in\mathcal{A}\cap\mathcal{H}_{0}}\frac{{{\widetilde{V}}}_{i}}{{\widetilde{R}}_{m_{a}}\lor 1}\right].$
@@
$\mathbb{E}\left[\sum_{i\in\mathcal{A}\cap\mathcal{H}_{0}}\frac{{{\widetilde{V}}}_{i}}{{\widetilde{R}}_{m_{a}}\lor 1}\right]\leq\mathbb{E}\left[\frac{|\mathcal{A}\cap\mathcal{H}_{0}|}{{\widetilde{R}}_{m_{a}}\lor 1}\right]\overset{(a)}{=}\mathbb{E}\left[\frac{m_{\mathcal{A}}}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$ (20)
@@
$\widetilde{W}_{i}\;=\;(\{S_{k+1},\dots,S_{n},S_{n+i}\}\;,\;(S_{n+j}:j\neq i,j\in H_{0})\;,\;(S_{n+j}:j\in H_{1})\,),$
@@
where the last equality is due to $\tau^{\prime}_{i}$ acting as a known constant inside that inner conditional. From Lemma 3, we know that $(n-k+1){\tilde{p}}_{i}$ is exactly the rank of $S_{n+i}$ among $\{S_{k+1},\dots,S_{n},S_{n+i}\}$ and ${\tilde{p}}_{i}$ is independent of $\widetilde{W}_{i}$. As a result, $(n-k+1)\,{\tilde{p}}_{i}$ is uniform over [1:nk+1][1:n-k+1], independent of $\tau^{\prime}_{i}$. Thus
@@
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{\lfloor\alpha(n-k+1)\,\tau^{\prime}_{i}/m\rfloor}{(n-k+1)\,\tau^{\prime}_{i}}\right]\leq m_{0}\cdot\mathbb{E}\left[\frac{\lfloor\alpha(n-k+1)\,\tau^{\prime}_{i}/m\rfloor}{(n-k+1)\,\tau^{\prime}_{i}}\right].$
The first term is thus bounded by $\alpha$.
Appendix D Direct Decision-based Attack Scheme

### v4

@@ III-C Surrogate Decision-based Attack Scheme
Motivated by our oracle setting, we consider the following practical scenario. The attacker does not have the training samples $\{Z_{j}\}_{j=1}^{n}$ and does not know the true labels of the test samples, but is allowed to query the label from the user who applies AdaDetect, with underlying machine learning algorithms unknown to the attacker. Such query access is a standard assumption in adversarial settings as mentioned in Section I-B, reflecting a realistic constraint on the attacker’s capability. Specifically, the attack has
@@
Query: The attacker can query the user (who owns all the training and test data and the AdaDetect algorithm) to obtain the labels for all the testing data $\{Z_{n+j}\}_{j=1}^{m}$. We denote these labels as $\{Y_{i}\}_{i=1}^{m}$ and refer to them as pseudo-labels, since they are not necessarily the true label in contrast to $Y_{i}^{*}$ in the oracle setting.
We propose a surrogate score function $g_{\text{surrogate}}(z)$, trained on the pseudo-labeled dataset $\mathcal{D}_{\text{surroage}}=\{(Z_{n+i},Y_{i})\}_{i=1}^{m}$. (Unlike the true labels available in our oracle setting, these are the labels assigned by AdaDetect on the entire test set.) We now describe the three steps of this attack method.
@@
Step 2: Surrogate score function training. Assuming AdaDetect has reasonable detection power, we form the pseudo-labeled dataset $\mathcal{D}_{\text{surrogate}}=\{(Z_{n+i},Y_{i})\}_{i=1}^{m}$ and train a surrogate score function
$g_{\text{surrogate}}(z)\leftarrow\mathrm{TrainScoreFunction}(\mathcal{D}_{\text{surrogate}}).$
@@
$\widetilde{Z}_{n+i}=f_{\text{attack-surroage}}(Z_{n+i}\,;\,g_{\text{surroage}}(z)).$ (20)
Remark 4.
One natural choice of $\mathcal{A}$ is to select the unrejected hypotheses with the smallest $p$-values (i.e., those closest to the rejection boundary). Let $(i_{1},i_{2},\ldots,i_{m-R})$ denote the indices of unrejected hypotheses ordered by their $p$-values: $\hat{p}_{i_{1}}\leq\hat{p}_{i_{2}}\leq\cdots\leq\hat{p}_{i_{m-R}}$. Then we define $\mathcal{A}=\{i_{1},i_{2},\ldots,i_{m_{\mathcal{A}}}\}$. This selects the $m_{\mathcal{A}}$ unrejected indices with the smallest $p$-values, targeting hypotheses that are close to $\hat{\tau}$ and making it an effective attack strategy as demonstrated in our experiments.
For this surrogate method, the main challenge in analyzing the FDR is because equation 7 no longer holds, which implies that $f_{\text{surrogate-attack}}$ does not satisfy Proposition 1. Specifically, recall that in the oracle setting, $g(z)$ is invariant to order of elements in $\{Z_{1},\dots,Z_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A}\}$ as they are all labeled as 0, which are the true labels known to the attacker. However, this fails to hold in the surrogate setting because (I) $g_{\text{surrogate}}(z)$ is trained only on the test data, and more importantly, (II) $g_{\text{surrogate}}(z)$ is trained on the pseudo-labels rather than the true labels, which breaks the invariance property.
It is important to note that our surrogate decision-based attack does not require information about the algorithm, making it a practical attack scheme. This point is also highlighted in our experiment section through mismatched setups, where the user and attacker adopt two different algorithms to learn the score function (see Experiment 3 and Experiment A.1 for details).
@@ IV Experiments
As explained in Section III-A, we will focus on two adversarial machine learning attacks: HSJA and Boundary attack. Recall that they minimize $\lVert z-{\tilde{z}}\rVert^{2}$ such that the label is flipped, which is inherited in all of our experimental evaluations. Each experiment is repeated $20$ times to calculate the mean and variance of the FDR and power. We denote the AdaDetect score function by $\tilde{s}(z)$ and the score function used for attack by g(z){g}(z) for oracle and $g_{\text{surrogate}}(z)$ for our surrogate method. The estimated upper bound according to equation 8 is computed by $\frac{m_{0}-m_{a}}{m}\alpha+m_{a}\cdot\frac{1}{20}\sum_{i=1}^{20}\left(\frac{1}{{\widetilde{R}}^{(i)}\lor 1}\right)$. This is computed for most of the experiments and listed in the tables. In Appendix D, we present the synthetic data generation and experiment results, along with one more real-world data experiment.
Real-world datasets:
  • $\bullet$
  • $\bullet$
    Credit Card: Transactions made by credit cards, some of which are fraudulent [56].
  • $\bullet$
    KDDCup99: Network connections that includes a variety of simulated intrusions [57].
  • $\bullet$
    Mammography: Features from mammograms and some with microcalcifications [58].
@@ IV-A Applying Surrogate Scheme on Adadetect
We evaluated the attack performance on three types of synthetic data distributions: independent Gaussian, non-Gaussian, and exchangeable Gaussian data, and four real-world datasets (Shuttle, Credit Card, KDDCup99, and Mammography) with diverse characteristics and application domains. All experiments use the following base parameters: training sample size $n=5000$, testing sample size $m=1000$, $k=4000$, true null data $m_{0}=900$, and significance level $\alpha=0.1$. Each experiment is repeated $20$ times to calculate the mean and variance of the FDR and power.
Experiment 1: Varying attack size with random forest (RF) models. We evaluate the impact of attack size on both oracle and surrogate attack performance using identical RF architectures for both score functions. The attack size $m_{a}$ sweeps between $10$, $30$, $50$ and $200$. We assess how each attack scheme performs under different attack scales with homogeneous RF configurations.
We only test HSJA in synthetic data experiments while both of HSJA and Boundary attack are evaluated in real-world data experiments. The result shows that the FDR for both attack schemes are under the theoretical upper bound, and oracle attack outperforms surrogate attack in non-gaussian case when the original power is low. Overall, both attack schemes successfully increase the FDR across varying attack sizes.
@@
Dataset Independent Gaussian Non-Gaussian Exchangeable Gaussian
original FDR $08\pm 0.03 0.0$$8\pm 0.04 0.0$$08\pm 0.04$
oracle ($m_{a}=10$) $11\pm 0.02 0.0$$08\pm 0.05$ $0.10\pm 0.01$
oracle ($m_{a}=30$) $0.24\pm 0.02$ $0.10\pm 0.06$ $0.28\pm 0.01$
oracle ($m_{a}=50$) $0.36\pm 0.02$ $0.40\pm 0.05$ $0.38\pm 0.02$
surrogate ($m_{a}=50$) $0.34\pm 0.02$ $0.20\pm 0.05$ $0.37\pm 0.02$
oracle ($m_{a}=200$) $0.67\pm 0.00$ $0.71\pm 0.01$ $0.69\pm 0.00$
surrogate ($m_{a}=200$) $0.67\pm 0.00$ $0.64\pm 0.01$ $0.67\pm 0.00$
TABLE II: Experiment 1: Power + RF
Dataset Independent Gaussian Non-Gaussian Exchangeable Gaussian
original power $0.96\pm 0.02$ $0.55\pm 0.06$ $1.00\pm 0.00$
oracle ($m_{a}=50$) $0.99\pm 0.01$ $0.87\pm 0.01$ $1.00\pm 0.00$
surrogate ($m_{a}=50$) $0.96\pm 0.02$ $0.65\pm 0.05$ $1.00\pm 0.00$
oracle ($m_{a}=200$) $0.99\pm 0.01$ $0.98\pm 0.01$ $1.00\pm 0.00$
surrogate ($m_{a}=200$) $0.96\pm 0.02$ $0.78\pm 0.05$ $1.00\pm 0.00$
Experiment 2: Real-world data with RF models. We evaluated both oracle and surrogate attack performance on the four real-world datasets using identical RF architectures for both score functions, with attack size $m_{a}=200$. This allows us to compare attack effectiveness across different real-world data characteristics. We compare the boundary attack with our default HSJA under both oracle and surrogate attack schemes, using identical RF architectures for both score functions with $m_{a}=200$. The result shows that our oracle and surrogate attack schemes and decision-based algorithms (HSJA and Boundary) can significantly increase the FDR in comparison to the original FDR. We also report the simulation with $m_{a}=10$ for the surrogate method with Boundary attack in Table IV.
TABLE III: Experiment 2: FDR + RF ($m_{a}=200$)
@@
TABLE IV: Experiment 2: FDR + RF ($m_{a}=10$)
Dataset Credit-card Shuttle KDD Mammography
original FDR $08\pm 0.03 0.0$$1\pm 0.00 0.0$$4\pm 0.02 0.0$$04\pm 0.08$
surrogate+bound. $10\pm 0.01 0.0$$9\pm 0.01 0.0$$9\pm 0.01 0.0$$07\pm 0.05$
estimated upper bound $0.18$ $0.18$ $0.18$ $0.20$
TABLE V: Experiment 2: Power + RF
@@
We make two important observations. Firstly, compared to the other three datasets, the original power on Mammography is relatively low ($\sim 0.48$). As a consequence, the surrogate method learns the score function from less accurate labels, leading to a larger gap in FDR between the oracle vs. the surrogate. Similar phenomena appear in the next real-world experiments, as well as the two in the Appendix D. Secondly, the power after attack often increases, since the attack targets data points near the decision boundary in practice, some of which belong to the alternative hypothesis and are thus more likely to be correctly rejected.
Experiment 3: Real-world data with mismatched configurations. We apply mismatched score function configurations (RF–NN) to the four real-world datasets with a fixed attack size of $m_{a}=200$, evaluating both the oracle and surrogate attack performance. This enables us to assess how different model combinations affect each attack type’s performance across various real-world scenarios. The results indicate that an attacker can employ a model different from the user’s and still inflate the FDR beyond the target level. Our experiments show that using the neural network configuration as the attacker’s model can substantially speed up the attack process. The results for RF-RF configuration is covered in Experiment $2$.
TABLE VI: Experiment 3: FDR + RF-NN
@@
TABLE VII: Experiment 4 with Monte Carlo adjustment
@@
TABLE VIII: Experiment 4 without Monte Carlo adjustment
@@
To demonstrate the generality of our attack framework, we evaluate the adversarial robustness of the one-class classifier-based method proposed by Bates et al. in [4]. This approach serves as a fundamental baseline. We utilize two real-world datasets and a synthetic Gaussian data distribution. The null (e.g., non-fraud) data is partitioned into a training set of size $n_{\text{train}}=4000$ and a calibration set of size $n_{\text{cal}}=1000$. The test set $\mathcal{D}_{\text{test}}$ consists of $m=2000$ samples, comprising $m_{0}=1800$ inliers (normal transactions) and $m_{1}=200$ outliers (e.g., fraudulent transactions). The $p$-values are computed using one-class classifier-based scheme, and the BH procedure is applied for FDR control.
@@
(2) Surrogate Training with Label Flipping: The attacker queries the one-class classifier-based scheme to obtain initial binary rejections. To train the surrogate model $g_{\text{surrogate}}(z)$ (a Multi-Layer Perceptron with one hidden layer of 100 units), we employ a heuristic label-flipping strategy: $75\%$ of the rejected samples with the highest $p$-values are relabeled as inliers in the training set. This encourages the surrogate to learn a more restrictive decision boundary, as only the most extreme outliers retain their original labels. Since HSJA and Boundary attack are decision-based attack, this surrogate model with tighter decision boundary for outliers will encourage those two decision-based attack to be more aggressive. We remark that a more conservative boundary (e.g., using a $50\%$ threshold or no flip at all) often fails here because it results in a loose rejection region, making it difficult for the attack to successfully push a sample labeled inlier into the real outlier space. The sample may successfully move into this loose boundary for outliers, but it’s still not extreme enough for one-class classifier-based scheme to notice. While the optimal threshold may vary across different datasets, this parameter serves as a tuning knob: increasing the flipping percentage makes the HSJA and Boundary attack more aggressive, facilitating more effective adversarial generation.
(3) Adversarial Generation: We utilize HSJA and Boundary attack on the surrogate model $g_{\text{surrogate}}(z)$ to generate adversarial perturbations for the target set $\mathcal{A}$. The attack is aiming to cross the decision boundary.
Experiment 4: Both real-world and synthetic data on Bates’s method. We evaluated the attack performance at significance levels $\alpha=0.1$ both with and without Monte Carlo adjustment (see Remark 1). The results quantify whether our surrogate approach succeed at attacking the standard one-class classifier-based scheme.
@@ V Discussion
Appendix A Proof of Theorem 1
With Lemma 1 and Lemma 2 in place, the proof of Theorem 1 largely follows from that of [7, Theorem 4.3]. We present it here for completeness.
@@ Appendix A Proof of Theorem 1
We can decompose $\text{FDR}^{*}_{\text{attack}}$ as follows,
$\hskip-1.00006pt\text{FDR}^{*}_{\text{attack}}\hskip-1.99997pt=\hskip-1.99997pt\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}\right]+\mathbb{E}\left[\sum_{i\in\mathcal{A}\cap\mathcal{H}_{0}}\frac{{{\widetilde{V}}}_{i}}{{\widetilde{R}}_{m_{a}}\lor 1}\right].$
@@
$\mathbb{E}\left[\sum_{i\in\mathcal{A}\cap\mathcal{H}_{0}}\frac{{{\widetilde{V}}}_{i}}{{\widetilde{R}}_{m_{a}}\lor 1}\right]\leq\mathbb{E}\left[\frac{|\mathcal{A}\cap\mathcal{H}_{0}|}{{\widetilde{R}}_{m_{a}}\lor 1}\right]\overset{(a)}{=}\mathbb{E}\left[\frac{m_{a}}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$ (22)
@@
$\widetilde{W}_{i}\;=\;$ $(\{S_{k+1},\dots,S_{n},S_{n+i}\},\;(S_{n+j}:j\neq i,j\in\mathcal{H}_{0}\setminus\mathcal{A})\;,\;$
$\hskip 120.00018pt(S_{n+j}:j\in\mathcal{H}_{1})\,\cup(S_{n+j}:j\neq i,j\in\mathcal{A})\,).$
@@
where the last equality is due to $\tau^{\prime}_{i}$ acting as a known constant conditional on $\widetilde{W}_{i}$. From Lemma 3, we know that $(n-k+1){\tilde{p}}_{i}$ is the rank of $S_{n+i}$ among $\{S_{k+1},\dots,S_{n},S_{n+i}\}$ and ${\tilde{p}}_{i}$ is independent of $\widetilde{W}_{i}$. As a result, $(n-k+1)\,{\tilde{p}}_{i}$ is uniform over [1:nk+1][1:n-k+1], independent of $\tau^{\prime}_{i}$. Thus
@@
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{\lfloor\alpha(n-k+1)\,\tau^{\prime}_{i}/m\rfloor}{(n-k+1)\,\tau^{\prime}_{i}}\right]\leq\frac{m_{0}-m_{a}}{m}\alpha.$ (26)
This completes the proof since equation 9 comes from upper-bounding equation 8 by $\alpha$.
Appendix B Direct Decision-based Attack Scheme